y=x+2
x and y intercept?

Answer by nyc_function(2741)

To find the x-intercept, let y = 0 and solve for x.

y = x + 2

0 = x + 2

0 - 2 = x

- 2 = x

To find the y-intercept, let x = 0 and solve for y.


y = x + 2

y = 0 + 2

y = 2

Done!